Rs) Plant Gain Gls)= st25t2 Closed -ap System Cpin (a): when the P-Cntrol Gain kls)=k changes the PI-trol S)-K(1+ SOIVE the chaVacteristic eruentin, Sketd the rat locus atont chouging C6) when the P-tl Cain kis)=e donges the ID-Cantrl Gain Kis) -Ks5)( Solve the Charmcteristic equation, Sketch the reot ous abnt chenying e:when the PCwntrol Gain 1S KIS)-K, Determine and Pole's psition damping Jatio o. 5 and Plot the S0 that the donain rts have Yoot latus e PD-Lontnl Gain is kis)-kst5), Detemine and Pole's fositian d)when Go that the donain rats have a damPi Yatis Pnd Plot the 2 Yoot ocus o-2 e): Nov we use the PGon: -k, PD Gain ps)-Kises), PI Gain: KiS)= k(It Compare the stedy-eror storterusng Frnal value theoren abont -20 ) usng le), Plot the P,PP.PI GAn on aph
Res Plant Cain Gls) St2st2 Cosed lop system Cpin Ca): when the P-Contl Gain e)= K changes the PI-tre-(1+ SoIve the chaacteristic euntin, etl the Yort acns alost Cb whan the P-Ctl Cain is)k dnges the D-Crtrl Gain Kes) = KUsts)(1) Solve the Chamcteristic equutim, Sketd the t us ant chrging Pole's forition (:when the-Cantrol Gain 1S K Determine and that the Jorain rts hare damping Jertio 5 and Plot the Yoot locus (3-e (d)har the PD-Contl Gain is kis) = KISHS), Deremne and Pele's fesition damPing Yatio nd Plot the Go that the doain ts have Yoot ocus e) Nov we use the PGin: -, PD Gain s)-Kises), PI Gan: kis)=K(It Cour Pave the Stenly -eror stmte using Final value theoren a bent k-20 ung Cel, Plot the P,PP.PI Gain on gafh
Answer #1

(1 ccs) HLS)= KCsto.2) 6C5) HCs) sCs+ 25+2) Yf 2st2 Characteristic cquation t 6es) HCS) o KCS+0.2) SCO+2St 2) SC2+2) tR(Sto.290 3 x180 270 Sum of Real partof poles- Sum of Real part of Centroid o Nenos P-z. O. 2) -1.8 -2+0.2 0.9 -1-1-0- RLE- Root loct80- tan ニーtan -90 t 180-180 +455139 1 96-34 Now Dd 83. 6S Gain kes) CS+5) (14 ) KC545) 42541) GCS)HCS) RCS+5)Csto2) Ges) HCSKCS45) CSto 2) Root locey plot. Ges)HCS) Sc25 2) No.of poles 3, No of zeot2 No of Root toey B1anc ber poles po ler> z enos NoBrealk point thom 14 GLS HCS- k . -+ 2s25) SY+5.25 + diff co.to s on both side and cuate to zero, e S2)(s45-251) - (242)+Sa)い K+2 Kt2 2 K2 4 ヒ=2 Now Ges) HCS)= +25+2, No-of Zeros 20 No 0f Poles 2 2 HP>z loccey Branche = p2 No.of Root No ot asymptodeGH 90 FIHfi+i C-1ti-iv 2j ab Po KCS+5) ECS+5) +20+2 +25+2+ヒい+5K KeS+5) 14 S4252 KCS+5) 3 TF = given) - 0.866 SY+ SC2+ヒ)ナ2+5kH97 +031 PO KC 4+ji) kC-1+1+5) j2 2GH 190 + tan 75.96 d104.03 104-03 PLB 9.2 9-5 2.55.96 Bocal point Yom 1+Gts HS) o -Ct25+2)Steady state eor (ess)- S ECS) e ct E Let Consides Res uct) Cinit step in put KIS)EK, ヒニ20 ECS) Ces) GLS) 425+2 . RCsD Ets) 1Ct+ Kis Ctar O. 2 -2) s8nt 4s RCs) ECS)= Csto.2)GCS) 1-+ S( S25+2) ess = SCSV+2 5+2) +k(s+0.2) 7 Css 20 C0.2) o+ KCO.2) ess o